# NGGAN: Noise Generation GAN Based on the Practical Measurement Dataset for Narrowband Powerline Communications

### v2

@@ I Introduction
To measure the similarity performances of GAN-based models, several performance metrics are adopted, including (1) maximum value, (2) mean value, (3) energy value, (4) standard deviation, (5) skewness, (6) kurtosis, (7) the number of peaks larger than the certain threshold value, (8) skewness of auto-correlation, (9) kurtosis of auto-correlation, (10) cyclic spectral density (CSD), (11) cyclic spectral coherence (CSC), (12) principal component analysis (PCA), and (13) Fréchet inception distance (FID).
@@ IV-A Performance Metrics
8) Skewness of auto-correlation: The auto-correlation sequence is defined as , where is defined as . By substituting the auto-correlation sequence and its mean into (8), the sample auto-correlation skewness value is obtained by
9) Kurtosis of auto-correlation: By substituting the auto-correlation sequence and its mean into (9), the auto-correlation peak value is obtained by
10) CSD and CSC: The auto-correlation of cyclo-stationary noise exhibits periodic characteristics and can be estimated from cyclic auto-correlation and cyclic power spectral density to identify the original cyclo-stationary noise [15]. Therefore, we employed CSD and CSC plots to individually examine how the intensity components of the noise changed at different frequencies and their correlations. The cyclic auto-correlation function (CAF), which is expressed as follows:
@@ V-A Dataset-1 (PSCGM-Generated)
Table IIIb (a) presents the statistics of cyclic auto-correlation coefficients exceeding 0.9 of noise generated by the four GAN-based models (15,000 samples) for cyclic frequencies of 122, 244, 366, 488, 610, and 732 Hz over the frequency range of 0 to 200 kHz. The table lists the average number of autocorrelation coefficients that exceed 0.9. Because all of the cyclic autocorrelation coefficients in Dataset-1 exceed 0.9, a higher percentage of autocorrelation coefficients of the generated data exceeding 0.9 indicates a higher similarity with Dataset-1. The error was calculated over the frequency range of 0-732 kHz. Statistical analysis showed that the NGGAN had the lowest accumulated error (33).
Table IIIb (b) presents the distribution of the maximum autocorrelation coefficients. If the number of maximum auto-correlation coefficients of the GAN-based models is closer to that of Dataset-1, it indicates a better learning performance in its cyclic spectral properties. Because the maximum auto-correlation coefficients of all GAN-based models were located at a cyclic frequency of 122 Hz, we focused on the maximum auto-correlation coefficients located at 244, 366, 488, 610, and 732 Hz. The NGGAN outperformed the other GAN-based models in four statistical quantities with errors of 5.
@@ V-B Dataset-2 (FRESH-Generated)
Table VIb (a) presents the statistics of cyclic auto-correlation coefficients exceeding 0.5 of noise generated by the four GAN-based models (15,000 samples) for the cyclic frequencies of 122, 244, 366, 488, 610, and 732 Hz. The table lists the average number of autocorrelation coefficients that exceed 0.5. By calculating the average number of correlation coefficients exceeding 0.5 over a frequency range of 0-200 kHz, we determined that the most important features were those associated with a cycling frequency of 122 Hz. The error was calculated over the frequency range of 0-122 kHz. Similarly, the four GAN-based models performed well in extracting features, as indicated by the error results.
Table VIb (b) presents the distributions of the maximum auto-correlation coefficients. For all the GAN-based models, the maximum autocorrelation coefficients were located at a cyclic frequency of 122 Hz. Therefore, we focused on the maximum autocorrelation coefficients at 244 Hz, 366 Hz, 488 Hz, 610 Hz, and 732 Hz. It is clear that the FD-SpecGAN outperformed all other GAN-based models, as indicated by a cumulative error of 7.
@@ V-C Dataset-3 (Experimentally Measured Data)
Table IXb (a) presents the statistics of the auto-correlation coefficients exceeding 0.3 of noise generated by the four GAN-based models generated by the four GAN-based models (15,000 samples) for cycling frequencies of 114, 228, 342, 456, 570, and 684 Hz. The error was calculated over the frequency range of 0-342 kHz. The table lists the average number of autocorrelation coefficients that exceed 0.3. Because all the cyclic autocorrelation coefficients in Dataset-3 exceed 0.3, the higher percentage of autocorrelation coefficients of the generated data exceeding 0.3 indicates a higher similarity with Dataset-3. Statistical analysis showed that the NGGAN had the lowest accumulated error (33). The simulation results revealed that the most important features in Dataset-3 were associated with cyclo-stationary frequencies of 114, 228, and 342 Hz. The FD-SpecGAN achieved the best performance with a cumulative error of 31, whereas the NGGAN achieved similar performance.
Table IXb (b) presents the distribution of maximum auto-correlation coefficients. The maximum autocorrelation coefficients of the four GAN-based models were located at a cyclic frequency of 114 Hz. Therefore, we focused on the maximum autocorrelation coefficients at 228, 342, 456, 570, and 684 Hz. The FD-SpecGAN outperformed the other GAN-based models in learning Dataset-3. However, the NGGAN achieved a performance similar to that of the FD-SpecGAN. The PL-SpecGAN had the poorest performance. As shown in Tables IXb (a) and (b), the NGGAN outperformed the DCGAN by 37 and 5, respectively.

### v3

@@ I Introduction
To measure the similarity performances of GAN-based models, several performance metrics are adopted, including (1) maximum value, (2) mean value, (3) energy value, (4) standard deviation, (5) skewness, (6) kurtosis, (7) the number of peaks larger than the certain threshold value, (8) skewness of autocorrelation, (9) kurtosis of autocorrelation, (10) cyclic spectral density (CSD), (11) cyclic spectral coherence (CSC), (12) principal component analysis (PCA), and (13) Fréchet inception distance (FID).
@@ IV-A Performance Metrics
8) Skewness of autocorrelation: The autocorrelation sequence is defined as , where is defined as . By substituting the autocorrelation sequence and its mean into (8), the sample autocorrelation skewness value is obtained by
9) Kurtosis of autocorrelation: By substituting the autocorrelation sequence and its mean into (9), the autocorrelation peak value is obtained by
10) CSD and CSC: The autocorrelation of cyclo-stationary noise exhibits periodic characteristics and can be estimated from cyclic autocorrelation and cyclic power spectral density to identify the original cyclo-stationary noise [15]. Therefore, we employed CSD and CSC plots to individually examine how the intensity components of the noise changed at different frequencies and their correlations. The cyclic autocorrelation function (CAF), which is expressed as follows:
@@ V-A Dataset-1 (PSCGM-Generated)
Table IIIb (a) presents the statistics of cyclic autocorrelation coefficients exceeding 0.9 of noise generated by the four GAN-based models (15,000 samples) for cyclic frequencies of 122, 244, 366, 488, 610, and 732 Hz over the frequency range of 0 to 200 kHz. The table lists the average number of autocorrelation coefficients that exceed 0.9. Because all of the cyclic autocorrelation coefficients in Dataset-1 exceed 0.9, a higher percentage of autocorrelation coefficients of the generated data exceeding 0.9 indicates a higher similarity with Dataset-1. The error was calculated over the frequency range of 0-732 kHz. Statistical analysis showed that the NGGAN had the lowest accumulated error (33).
Table IIIb (b) presents the distribution of the maximum autocorrelation coefficients. If the number of maximum autocorrelation coefficients of the GAN-based models is closer to that of Dataset-1, it indicates a better learning performance in its cyclic spectral properties. Because the maximum autocorrelation coefficients of all GAN-based models were located at a cyclic frequency of 122 Hz, we focused on the maximum autocorrelation coefficients located at 244, 366, 488, 610, and 732 Hz. The NGGAN outperformed the other GAN-based models in four statistical quantities with errors of 5.
@@ V-B Dataset-2 (FRESH-Generated)
Table VIb (a) presents the statistics of cyclic autocorrelation coefficients exceeding 0.5 of noise generated by the four GAN-based models (15,000 samples) for the cyclic frequencies of 122, 244, 366, 488, 610, and 732 Hz. The table lists the average number of autocorrelation coefficients that exceed 0.5. By calculating the average number of correlation coefficients exceeding 0.5 over a frequency range of 0-200 kHz, we determined that the most important features were those associated with a cycling frequency of 122 Hz. The error was calculated over the frequency range of 0-122 kHz. Similarly, the four GAN-based models performed well in extracting features, as indicated by the error results.
Table VIb (b) presents the distributions of the maximum autocorrelation coefficients. For all the GAN-based models, the maximum autocorrelation coefficients were located at a cyclic frequency of 122 Hz. Therefore, we focused on the maximum autocorrelation coefficients at 244 Hz, 366 Hz, 488 Hz, 610 Hz, and 732 Hz. It is clear that the FD-SpecGAN outperformed all other GAN-based models, as indicated by a cumulative error of 7.
@@ V-C Dataset-3 (Experimentally Measured Data)
Table IXb (a) presents the statistics of the autocorrelation coefficients exceeding 0.3 of noise generated by the four GAN-based models generated by the four GAN-based models (15,000 samples) for cycling frequencies of 114, 228, 342, 456, 570, and 684 Hz. The error was calculated over the frequency range of 0-342 kHz. The table lists the average number of autocorrelation coefficients that exceed 0.3. Because all the cyclic autocorrelation coefficients in Dataset-3 exceed 0.3, the higher percentage of autocorrelation coefficients of the generated data exceeding 0.3 indicates a higher similarity with Dataset-3. Statistical analysis showed that the NGGAN had the lowest accumulated error (33). The simulation results revealed that the most important features in Dataset-3 were associated with cyclo-stationary frequencies of 114, 228, and 342 Hz. The FD-SpecGAN achieved the best performance with a cumulative error of 31, whereas the NGGAN achieved similar performance.
Table IXb (b) presents the distribution of maximum autocorrelation coefficients. The maximum autocorrelation coefficients of the four GAN-based models were located at a cyclic frequency of 114 Hz. Therefore, we focused on the maximum autocorrelation coefficients at 228, 342, 456, 570, and 684 Hz. The FD-SpecGAN outperformed the other GAN-based models in learning Dataset-3. However, the NGGAN achieved a performance similar to that of the FD-SpecGAN. The PL-SpecGAN had the poorest performance. As shown in Tables IXb (a) and (b), the NGGAN outperformed the DCGAN by 37 and 5, respectively.